The BRS Corporation makes collections on sales according to the following schedule:

40% in month of sale
56% in month following sale
4% in second month following sale

The following sales have been budgeted:

Sales
April $ 150,000
May $ 170,000
June $ 160,000

Budgeted cash collections in June would be:

Multiple Choice:

A) $165,200

B) $159,200

C) $160,600

D) $160,000
Answer #1

Budgeted cash collection in the month of June:

Tje following cash receipts are expecting in the month of June:

1. 40% in month of sale:

Here June month budgeted sale is $160,000.

40 % of the budgeted sale of June = $160,000 × 40/100 = $64,000

2. 56% in month following sale:

That means, 56% of the sales amount of month May will receive in the following month, ie, June. May month's budgeted sales is $170,000.

56% of the budgeted sale of May = $170,000 × 56/100 = $95,200

3. 4% in second month following sale:

That means, 4% of the sales amount of month April will receive in the second month following sale, ie, June. April month's budgeted sales is $150,000.

4% of the budgeted sale of April = $150,000 × 4/100 = $6,000

Total.cash collections = $64,000 + $95,200 + $6,000 = $165,200.

Correct answer is a) $165,200.